Paula mixed 2 cups of sunflower seeds and 3 cups of raisins to make a snack for a hike. She figured that the mixture would provide her with 2900 Calories of food energy. Find the number of Calories per cup of raisins if it is 400 less than the number of Calories per cup of sunflower seeds.
Answer by mananth(16946)

Paula mixed 2 cups of sunflower seeds
3 cups of raisins
mixture 2900 Calories energy.
raisins is 400 less than the number of Calories per cup of sunflower seeds.
let calories/cup in sunflower be x
calories /cup in raisins be y
..
x-y=400.......................1
..
2x+3y=2900....................2
multiply equation 1 by 3 and add to equation 2
3x-3y=1200
2x+3y=2900
ADD
5x=4100
divide by x
x= 4100/5
x=820 calories /cup in sun flower.
raisins contain 400 less. 820-400 = 420 calories